THZ SPECTROSCOPY OF ACETALDEHYDE AND SEARCH OF ^{13}C SPECIES IN ORION

\underline{L. MARGULÈS}, R. A. MOTIYENKO, Laboratoire PhLAM, CNRS UMR 8523, Université de Lille 1, 59655 Villeneuve d’Ascq Cedex, France; V. V. ILYUSHIN, Institute of Radio Astronomy of NASU, Chervonopraporna Str., 4, 61002 Kharkov, Ukraine; B. TERCERO, J. CERNICHARO, Centro de Astrobiología (CSIC-INTA). Laboratory of Molecular Astrophysics. Department of Astrophysics. Ctra de Ajalvir,Km 4, 28850 Torrejón de Ardoz, Madrid, Spain; and J.-C. GUILLEMIN, Sciences Chimiques de Rennes, UMR 6226 CNRS-ENSCR, Avenue du Général Leclerc, CS 50837, 35708 Rennes Cedex 7, France.

Acetaldehyde (CH_{3}CHO) is one of the high priority complex organic molecules for the astrophysical community. There is a lack of data concerning the ^{13}C species since the measurements are limited to 40 GHz^{a} up to date. This molecule displays a large amplitude motion: the hindered rotation of the methyl group with respect to the rest of the molecule. The analysis is performed with RAM36 code^{b} which used the Rho Axis Method. Last year we presented the analysis of the millimeterwave spectra of the ^{13}CH_{3}CHO species^{c}. We extended the analysis to the THz range of the vibrational ground state for both species. We are also analyzing the first torsional state (≈140 cm^{−1}) for two reasons: first, this permits to remove correlation between parameters. Second, this state contribute to the partition function even at ISM temperature (100–150 K) since there is an influence on the column density determined in case of detection. The searches of these isotopomers are in progress in ORION. This work was supported by the CNES and the Action sur Projets de l’INSU, PCMI. This work was also done under the ANR-08-BLAN-0054.

aKilb, R.W.; Lin, C.C.; and Wilson, E.B.J. Chem. Phys. 26, (1957) 1695

bIlyushin, V.V. et al;J. Mol. Spectrosc. 259, (2010) 26

cMargulès, L. et al; FA07, 66th International Symposium on Molecular Spectroscopy (2011)